OpenStudy (anonymous):

-4√x+9 (end of sqrt) = 20

OpenStudy (anonymous):

\[-4\sqrt{x+9} = 20\]

OpenStudy (anonymous):

is this the question??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok so squaring bothsides we get 16(x+9) = 400 x+9 = 400/16 x+9 = 25 x = 25 - 9 x = 16 --------------> required solution

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

welcome